# dUltra: Ultra-Fast Diffusion Language Models via Reinforcement Learning

### v1

@@ Abstract
Masked diffusion language models (MDLMs) offer the potential for parallel token generation, but most open-source MDLMs decode fewer than 5 tokens per model forward pass even with sophisticated sampling strategies. As a result, their sampling speeds are often comparable to AR + speculative decoding schemes, limiting their advantage over mainstream autoregressive approaches. Existing distillation-based accelerators (dParallel, d3LLM) finetune MDLMs on trajectories generated by a base model, which can become off-policy during finetuning and restrict performance to the quality of the base model’s samples. We propose dUltra, an on-policy reinforcement learning framework based on Group Relative Policy Optimization (GRPO) that learns unmasking strategies for efficient parallel decoding. dUltra introduces an unmasking planner head that predicts per-token unmasking likelihoods under independent Bernoulli distributions. We jointly optimize the base diffusion LLM and the unmasking order planner using reward signals combining verifiable reward, distillation reward, and the number of unmasking steps. Across mathematical reasoning and code generation tasks, dUltra improves the accuracy–efficiency trade-off over state-of-the-art heuristic and distillation baselines, moving towards achieving “diffusion supremacy” over autoregressive models.
@@ 1 Introduction
The great success of diffusion models and their parallel generation nature makes them attractive for language generation. There has been an ongoing effort to transfer the diffusion paradigm to the natural language domain. Earlier attempts tried to sample in the token embedding space (NEURIPS2022_1be5bc25; gulrajani2023likelihoodbaseddiffusionlanguagemodels; gong2023diffuseqsequencesequencetext) or the latent space of a language autoencoder (lovelace2023latentdiffusionlanguagegeneration) using continuous diffusion. In a parallel line of work, discrete diffusion in the vocabulary space is studied in (austin2023structureddenoisingdiffusionmodels; campbell2022continuous; meng2022concrete; he2022diffusionbertimprovinggenerativemasked; wang2019bertmouthspeakbert). Without the need for an invertible token-wise embedding function, discrete diffusion language models generally perform better than continuous diffusion language models when operating in a discrete state space with a countable number of states (lou_discrete_2024; shi_simplified_2024; sahoo_simple_2024). Moreover, it is observed that the best performance is achieved with an absorbing source distribution, which further restricts possible sampling paths (lou_discrete_2024; austin2023structureddenoisingdiffusionmodels). The diffusion models with an absorbing source distribution are also called masked diffusion language models (MDLMs) because of the introduction of a masked token. Moreover, ou2025absorbingdiscretediffusionsecretly and sahoo_simple_2024 have shown that the optimal solution that minimizes negative ELBO is independent of the time variable, thus promoting time-agnostic parameterization. Recent works have scaled time-agnostic MDLMs to show that their performance can rival that of autoregressive (AR) models with similar parameter size (nie2025largelanguagediffusionmodels; dream2025).
@@
However, even these best open-source masked diffusion language models (MDLMs) (nie2025largelanguagediffusionmodels; dream2025) face the same slow sampling problem as continuous diffusion models. Apart from system-level optimization, such as approximate KV Cache (wu2025fastdllmtrainingfreeaccelerationdiffusion; hu2025acceleratingdiffusionlanguagemodel), attempts to accelerate MDLMs have focused on designing confidence/entropy-aware sampling methods to unmask tokens in parallel (ben-hamu_accelerated_2025; wu2025fastdllmtrainingfreeaccelerationdiffusion). These deterministic parallel decoding strategies generally favor tokens with high confidence or low entropy within a single unmasking step, leading to 3-5x faster sampling speed. Despite this decent speed-up, the number of tokens decoded in parallel within a single denoising step is still low ($\sim$4.5 tokens per step on GSM8K and $\sim$3.3 tokens per step on MATH500) (wu2025fastdllmtrainingfreeaccelerationdiffusion). Meanwhile, commercial diffusion language models (gemini-diffusion; labs2025mercury) operate at a throughput of $\sim$1000 tokens/s, significantly surpassing the state-of-the-art throughput of autoregressive models, which typically range from 100-300 tokens/s (labs2025mercury). Therefore, current open-source MDLMs have yet to exploit the parallel generation potential of diffusion language models to their fullest and have not achieved "diffusion supremacy" (analogous to quantum supremacy) over autoregressive models.
@@
A fundamental observation is that when tokens are generated in parallel, MDLMs assume conditional independence between them given the masked input sequence (see also Section˜2.2.2) due to the curse of dimensionality (dllm_ratio_blog). If the tokens are actually dependent and ambiguous [^1] An ambiguous token means that multiple words in the vocabulary are predicted to have similar and high probability., the actual distribution sampled from will have unwanted modes, and the model may generate nonsensical text (Figure˜1, first row, middle panel). Since acceleration methods aim to increase parallelism by decoding more tokens per step, they inevitably unmask some dependent tokens together, violating the independence assumption. Therefore, given the current paradigm of parallel decoding [^2] Unless curse of dimensionality is solved, any method that accelerates sampling is implementing some form of mode filtering, i.e. the accelerated sampling procedure will try to sample from a single mode that produces sensical text out of all modes that a slower model can sample from. Figure˜1 illustrates this behavior.
@@
Recent work (d3llm; chen2025dparallellearnableparalleldecoding) on accelerating diffusion LLM sampling with offline distillation demonstrates this principle as well. Specifically, d3LLM (d3llm) distills from the sampling trajectory produced by the teacher model, effectively teaching the student model to sample from one of the high probability teacher modes at each sampling step. Similarly, dParallel (chen2025dparallellearnableparalleldecoding) proposes certainty-forcing distillation that finetunes the model to be more confident on self-generated trajectories, which are fixed per prompt. This effectively encourages the model to sample from one of the high probability teacher modes marginalized by the prompts. However, this self-distillation approach faces two shortcomings. First of all, the quality of training data cannot exceed the capability of the base model. In other words, the training data only includes greedily sampled trajectories, limiting exploration and performance of the distilled model. Secondly, the self-generated trajectory can be off-policy during distillation, and the student learns in contexts more frequently encountered by the teacher model rather than by itself. Therefore, we want a method that explores the teacher model modes and trains on samples generated by the student model. On-policy reinforcement learning naturally addresses both limitations: by training on trajectories generated by the current policy, it avoids the distribution mismatch inherent in offline distillation, while the exploration encouraged by RL allows discovering better unmasking strategies beyond those demonstrated by the base model.
@@
  • We develop a GRPO-based training framework that jointly optimizes the diffusion model and unmasking order planner using multiple reward signals including verifiable reward, distillation reward, and unmasking efficiency.
  • We propose advantage clipping to enable effective RL training of the unmasking planner head.
  • We demonstrate our approach’s strong performance on mathematical reasoning (GSM8K, MATH500) and code generation (HumanEval, MBPP) tasks.
@@ 2.2.1 Single token case
Continuous-time MDLMs (shi_simplified_2024; sahoo_simple_2024) can be formulated with a forward noising process and the corresponding reversal. For the forward noising process, we have for $t,s\in[0,1]$ and $t>s$,
@@
Similar to DDPM (ho_denoising_2020), the parameterization of MDLM $\bm{\mu}_{\theta}$ aims to predict the clean sample $x_{0}$:
@@ 2.2.3 Likelihood bound
For MDLM with carry-over unmasking (sahoo_simple_2024; shi_simplified_2024; nie2025largelanguagediffusionmodels; dream2025), the NELBO loss can be simplified as
@@ 2.2.4 Connection to any-order autoregressive modeling
Any-order autoregressive models (AO-ARMs) (uria2014deeptractabledensityestimator; hoogeboom2021argmaxflowsmultinomialdiffusion; shih2022traininginferenceanyorderautoregressive) evaluate the likelihood of a sequence $\bm{x}$ by marginalizing over all possible generation orders. More specifically, the expected negative log-likelihood minimized by AO-ARMs during training is
@@
where $U(S_{\left|\bm{x}\right|})$ is the uniform distribution over all permutations over the initial mask tokens, and each permutation uniquely specifies a sampling order. Works have shown that this loss is equivalent to the NELBO loss of MDLMs with carry-over unmasking (Equation˜2.6) under mild conditions (ou2025absorbingdiscretediffusionsecretly; kim_train_2025). Therefore, MDLMs with carry-over unmasking can be interpreted as AO-ARMs that sample generation orders from a uniform distribution over all possible permutations.
@@ 3.3 Initializing on-policy scheduler with heuristic methods
More specifically, we finetune the unmasking planner head using supervised learning to mimic a confidence-based sampling strategy. We train the planner to predict token selection decisions made by the confidence-aware parallel decoding method proposed in Fast-dLLM (wu2025fastdllmtrainingfreeaccelerationdiffusion). This method selects highly confident tokens to unmask in each denoising step. This avoids the pathological case described in Figure˜1, and provides a strong baseline for our learnable planner.
@@ 3.4 On-policy distillation reward
We use domain-specific teacher models to provide high-quality supervision signals: Qwen2.5-Math-7B-Instruct (qwen2.5math) for mathematical reasoning and Qwen2.5-Coder-7B-Instruct (qwen2.5coder) for code generation. Given a prompt $\bm{q}$ and a completion $\bm{c}$ generated by the student model during rollout, we compute the on-policy distillation reward as:
@@
where $|\bm{c}|$ is the length of the completion in tokens, $p_{\text{student}}(\bm{c}\mid\bm{q})$ is given by Equation˜3.4, and $\beta$ controls the weight of the entropy regularization term. When $\beta=1$, maximizing ${\mathbb{E}}_{\bm{c}\sim p_{\text{student}}}[r_{\text{distill}}]$ corresponds to minimizing a scaled reverse KL divergence $\frac{1}{|\bm{c}|}D_{\text{KL}}(p_{\text{student}}\|p_{\text{teacher}})$. Reverse KL divergence is known to encourage mode-seeking behavior (minka2005divergence): when the student places probability mass where the teacher model has low probability, the term $\log(p_{\text{student}}/p_{\text{teacher}})$ becomes large and heavily penalizes the divergence. We set $\beta=0$ in our experiments, simplifying the reward to average teacher log-likelihood, while still encouraging mode-seeking behavior: the student learns to generate completions that the teacher assigns high probability, focusing on high-quality modes verified by the teacher model rather than exploring low-probability regions that may lead to incoherent text. Empirically, we did not find significant differences in performance when varying $\beta$.
@@ 3.5 GRPO Training Framework
We train the diffusion LLM and unmasking planner head using Group Relative Policy Optimization (GRPO), a policy-gradient method that operates on groups of trajectories per prompt and uses group-relative advantages as baselines. We use on-policy updates (num_iterations=1=1) as off-policy updates provide no computational benefit: unlike autoregressive LLMs that evaluate likelihoods in a single forward pass, each likelihood evaluation using Equation˜3.2 requires a complete denoising process. For each group, we sample a single prompt $\bm{q}$ and generate $G$ trajectories with denoised sequence at step $k$ of $g$-th trajectory denoted as $\bm{o}^{g,k}$. The group-relative advantages of these trajectories are calculated as $A^{g}=r^{g}-\text{mean}(r^{1:G})$. Note that we do not normalize the advantage to avoid introducing any bias into the policy gradient estimation (liu2025understanding). Then the GRPO objective reduces to the REINFORCE-style loss
@@
In summary, the complete training algorithm can be found in Appendix˜B.
@@ 4 Results
We evaluate dUltra on mathematical reasoning and code generation benchmarks to demonstrate its ability to achieve competitive performance with significantly reduced computational cost. All of our experiments use LLaDA-8B-Instruct as the base model. We compare against state-of-the-art confidence-based heuristics, such as Fast-dLLM (wu2025fastdllmtrainingfreeaccelerationdiffusion), and state-of-the-art off-policy distillation-based methods, such as d3LLM (d3llm) and dParallel (chen2025dparallellearnableparalleldecoding). We report both task performance and the average number of function evaluations (NFE), i.e., the number of denoising steps during generation.
@@ 4.2.1 Results with Small Block Size
Fast-dLLM 77.56 57.26 37.20 77.12 33.54 62.65 37.80 50.04
dParallel 77.10 28.47 32.60 35.21 31.71 30.68 36.20 22.77
d3LLM 72.50 25.79 31.80 32.04 37.20 33.41 33.20 32.26
dUltra-math (ours) 81.29 21.21 35.52 33.20 - - - -
   w/o planner 75.44 15.31 34.40 24.67 - - - -
dUltra-coding (ours) 81.52 23.54 35.64 33.39 35.85 24.29 37.04 19.78
   w/o planner 75.89 15.27 32.60 22.49 23.78 16.18 20.40 12.18
Table 1: Main results comparing dUltra with baseline methods across mathematical reasoning and code generation tasks. The generation length is 256 and the block size is 32 for all results in this table. We report accuracy and NFE (number of function evaluations) for each task. Bold indicates the best result and indicates the second-best result for each column. The “w/o planner” rows show results using Fast-dLLM sampling with the base model’s logits (without the learned unmasking planner head), serving as an ablation baseline. dUltra-math is trained on GSM8K training dataset while dUltra-coding uses dUltra-math as the base checkpoint and is trained on the APPS dataset.

### v2

@@ Abstract
Masked diffusion language models (MDLMs) offer the potential for parallel token generation, but most open-source MDLMs decode fewer than 5 tokens per model forward pass even with sophisticated sampling strategies, limiting their parallel generation potential. Existing acceleration methods either rely on fixed confidence-based heuristics or use distillation-based approaches that finetune MDLMs on trajectories generated by a base model, which can become off-policy during finetuning and restrict performance to the quality of the base model’s samples. We propose dUltra, an on-policy reinforcement learning framework based on Group Relative Policy Optimization (GRPO) that learns unmasking strategies for efficient parallel decoding. dUltra introduces an unmasking planner head that predicts per-token unmasking likelihoods under independent Bernoulli distributions. We jointly optimize the base diffusion LLM and the unmasking order planner using reward signals combining verifiable reward, distillation reward, and the number of unmasking steps. Across mathematical reasoning and code generation tasks, dUltra achieves superior accuracy-efficiency trade-offs compared to state-of-the-art heuristic (Fast-dLLM) and distillation baselines (d3LLM, dParallel), demonstrating that learned unmasking trajectories through on-policy RL enable better exploitation of parallel generation in MDLMs. Code and checkpoints are released here.
@@ 1 Introduction
The great success of diffusion models and their parallel generation nature makes them attractive for language generation. There has been an ongoing effort to transfer the diffusion paradigm to the natural language domain. Earlier attempts tried to sample in the token embedding space (Li et al., 2022; Gulrajani and Hashimoto, 2023; Gong et al., 2023) or the latent space of a language autoencoder (Lovelace et al., 2023) using continuous diffusion. In a parallel line of work, discrete diffusion in the vocabulary space is studied in (Austin et al., 2021; Campbell et al., 2022; Meng et al., 2022; He et al., 2022; Wang and Cho, 2019). Without the need for an invertible token-wise embedding function, discrete diffusion language models generally perform better than continuous diffusion language models when operating in a discrete state space with a countable number of states (Lou et al., 2023; Shi et al., 2024; Sahoo et al., 2024). Moreover, it is observed that the best performance is achieved with an absorbing source distribution, which further restricts possible sampling paths (Lou et al., 2023; Austin et al., 2021). The diffusion models with an absorbing source distribution are also called masked diffusion language models (MDLMs) because of the introduction of a masked token. Moreover, Ou et al. (2025) and Sahoo et al. (2024) have shown that the optimal solution that minimizes negative ELBO is independent of the time variable, thus promoting time-agnostic parameterization. Recent works have scaled time-agnostic MDLMs to show that their performance can rival that of autoregressive (AR) models with similar parameter size (Nie et al., 2025; Ye et al., 2025).
@@
However, even these best open-source masked diffusion language models (MDLMs) (Nie et al., 2025; Ye et al., 2025) face the same slow sampling problem as continuous diffusion models. Apart from system-level optimization, such as approximate KV Cache (Wu et al., 2025; Hu et al., 2025), attempts to accelerate MDLMs have focused on designing confidence/entropy-aware sampling methods to unmask tokens in parallel (Ben-Hamu et al., 2025; Wu et al., 2025). These deterministic parallel decoding strategies generally favor tokens with high confidence or low entropy within a single unmasking step, leading to 3-5x faster sampling speed. Despite this decent speed-up, the number of tokens decoded in parallel within a single denoising step is still low ($\sim$4.5 tokens per step on GSM8K and $\sim$3.3 tokens per step on MATH500) (Wu et al., 2025). Meanwhile, commercial diffusion language models (Google DeepMind, 2025; Khanna et al., 2025) operate at a throughput of $\sim$1000 tokens/s, significantly surpassing the state-of-the-art throughput of autoregressive models, which typically range from 100-300 tokens/s (Khanna et al., 2025). Therefore, current open-source MDLMs have yet to exploit the parallel generation potential of diffusion language models to their fullest and have not achieved "diffusion supremacy" (analogous to quantum supremacy) over autoregressive models.
@@
A fundamental observation is that when tokens are generated in parallel, MDLMs assume conditional independence between them given the masked input sequence (see also Section˜2.2.2) due to the curse of dimensionality (Aaron, 2024). If the tokens are actually dependent and ambiguous [^1] An ambiguous token means that multiple words in the vocabulary are predicted to have similar and high probability., the actual distribution sampled from will have unwanted modes, and the model may generate nonsensical text (Figure˜1, first row, middle panel). Since acceleration methods aim to increase parallelism by decoding more tokens per step, they inevitably unmask some dependent tokens together, violating the independence assumption. Therefore, given the current paradigm of parallel decoding [^2] Unless curse of dimensionality is solved, any method that accelerates sampling is implementing some form of mode filtering, i.e. the accelerated sampling procedure will try to sample from a single mode that produces sensical text out of all modes that a slower model can sample from. Figure˜1 illustrates this behavior.
@@
Recent work (Qian et al., 2025; Chen et al., 2025) on accelerating diffusion LLM sampling with offline distillation demonstrates this principle as well. Specifically, d3LLM (Qian et al., 2025) distills from the sampling trajectory produced by the teacher model, effectively teaching the student model to sample from one of the high probability teacher modes at each sampling step. Similarly, dParallel (Chen et al., 2025) proposes certainty-forcing distillation that finetunes the model to be more confident on self-generated trajectories, which are fixed per prompt. This effectively encourages the model to sample from one of the high probability teacher modes marginalized by the prompts. However, this self-distillation approach faces two shortcomings. First of all, the quality of training data cannot exceed the capability of the base model. In other words, the training data only includes greedily sampled trajectories, limiting exploration and performance of the distilled model. Secondly, the self-generated trajectory can be off-policy during distillation, and the student learns in contexts more frequently encountered by the teacher model rather than by itself. Therefore, we want a method that explores the teacher model modes and trains on samples generated by the student model. On-policy reinforcement learning naturally addresses both limitations: by training on trajectories generated by the current policy, it avoids the distribution mismatch inherent in offline distillation, while the exploration encouraged by RL allows discovering better unmasking strategies beyond those demonstrated by the base model.
@@
  • We develop a GRPO-based on-policy distillation framework that jointly optimizes the diffusion model and unmasking order planner using both verifiable reward and distillation reward. We also propose advantage clipping to enable effective RL training of the unmasking planner head.
  • We show state-of-the-art performance in terms of both token-per-forward (TPF) and number of function evaluations (NFE) on mathematical reasoning (GSM8K, MATH500) and code generation (HumanEval, MBPP) tasks.
@@ 2.2.1 Single token case
Continuous-time MDLMs (Shi et al., 2024; Sahoo et al., 2024) can be formulated with a forward noising process and the corresponding reversal. For the forward noising process, we have for $t,s\in[0,1]$ and $t>s$,
@@
Similar to DDPM (Ho et al., 2020), the parameterization of MDLM $\bm{\mu}_{\theta}$ aims to predict the clean sample $x_{0}$:
@@ 2.2.3 Likelihood bound
For MDLM with carry-over unmasking (Sahoo et al., 2024; Shi et al., 2024; Nie et al., 2025; Ye et al., 2025), the NELBO loss can be simplified as
@@ 2.2.4 Connection to any-order autoregressive modeling
Any-order autoregressive models (AO-ARMs) (Uria et al., 2014; Hoogeboom et al., 2021; Shih et al., 2022) evaluate the likelihood of a sequence $\bm{x}$ by marginalizing over all possible generation orders. More specifically, the expected negative log-likelihood minimized by AO-ARMs during training is
@@
where $U(S_{\left|\bm{x}\right|})$ is the uniform distribution over all permutations over the initial mask tokens, and each permutation uniquely specifies a sampling order. Works have shown that this loss is equivalent to the NELBO loss of MDLMs with carry-over unmasking (Equation˜2.6) under mild conditions (Ou et al., 2025; Kim et al., 2025). Therefore, MDLMs with carry-over unmasking can be interpreted as AO-ARMs that sample generation orders from a uniform distribution over all possible permutations.
@@ 3.3 Initializing on-policy scheduler with heuristic methods
More specifically, we finetune the unmasking planner head using supervised learning to mimic a confidence-based sampling strategy. We train the planner to predict token selection decisions made by the confidence-aware parallel decoding method proposed in Fast-dLLM (Wu et al., 2025). This method selects highly confident tokens to unmask in each denoising step. This avoids the pathological case described in Figure˜1, and provides a strong baseline for our learnable planner.
@@ 3.4 On-policy distillation reward
We use domain-specific teacher models to provide high-quality supervision signals: Qwen2.5-Math-7B-Instruct (Yang et al., 2024) for mathematical reasoning and Qwen2.5-Coder-7B-Instruct (Hui et al., 2024) for code generation. Given a prompt $\bm{q}$ and a completion $\bm{c}$ generated by the student model during rollout, we compute the on-policy distillation reward as:
@@
where $|\bm{c}|$ is the length of the completion in tokens, $p_{\text{student}}(\bm{c}\mid\bm{q})$ is given by Equation˜3.4, and $\beta$ controls the weight of the entropy regularization term. When $\beta=1$, maximizing ${\mathbb{E}}_{\bm{c}\sim p_{\text{student}}}[r_{\text{distill}}]$ corresponds to minimizing a scaled reverse KL divergence $\frac{1}{|\bm{c}|}D_{\text{KL}}(p_{\text{student}}\|p_{\text{teacher}})$. Reverse KL divergence is known to encourage mode-seeking behavior (Minka and others, 2005): when the student places probability mass where the teacher model has low probability, the term $\log(p_{\text{student}}/p_{\text{teacher}})$ becomes large and heavily penalizes the divergence. We set $\beta=0$ in our experiments, simplifying the reward to average teacher log-likelihood, while still encouraging mode-seeking behavior: the student learns to generate completions that the teacher assigns high probability, focusing on high-quality modes verified by the teacher model rather than exploring low-probability regions that may lead to incoherent text. Empirically, we did not find significant differences in performance when varying $\beta$.
@@ 3.5 GRPO Training Framework
We train the diffusion LLM and unmasking planner head using Group Relative Policy Optimization (GRPO), a policy-gradient method that operates on groups of trajectories per prompt and uses group-relative advantages as baselines. We use on-policy updates (num_iterations=1=1) as off-policy updates provide no computational benefit: unlike autoregressive LLMs that evaluate likelihoods in a single forward pass, each likelihood evaluation using Equation˜3.2 requires a complete denoising process. For each group, we sample a single prompt $\bm{q}$ and generate $G$ trajectories with denoised sequence at step $k$ of $g$-th trajectory denoted as $\bm{o}^{g,k}$. The group-relative advantages of these trajectories are calculated as $A^{g}=r^{g}-\text{mean}(r^{1:G})$. Note that we do not normalize the advantage to avoid introducing any bias into the policy gradient estimation (Liu et al., 2025b). Then the GRPO objective reduces to the REINFORCE-style loss
@@
In summary, the complete training algorithm can be found in Appendix B.
@@ 4 Results
We evaluate dUltra on mathematical reasoning and code generation benchmarks to demonstrate its ability to achieve competitive performance with significantly reduced computational cost. All of our experiments use LLaDA-8B-Instruct as the base model. We compare against state-of-the-art confidence-based heuristics, such as Fast-dLLM (Wu et al., 2025), and state-of-the-art off-policy distillation-based methods, such as d3LLM (Qian et al., 2025) and dParallel (Chen et al., 2025). We report both task performance and the average number of function evaluations (NFE), i.e., the number of denoising steps during generation.
@@ 4.2.1 Results with Small Block Size
Fast-dLLM 77.56 57.26 (4.33) 37.20 77.12 (3.22) 33.54 62.65 (3.94) 37.80 50.04 (4.16)
dParallel 77.10 28.47 (7.17) 31.80 35.21 (5.32) 31.71 30.68 (5.87) 36.20 22.77 (6.30)
d3LLM 72.50 25.79 (8.39) 32.60 32.04 (6.81) 37.20 33.41 (6.86) 33.20 32.26 (6.90)
dUltra-math (ours) 81.29 21.21 (9.717) 35.52 33.20 (7.03) - - - -
   w/o planner 75.44 15.31 (13.39) 34.00 24.67 (9.45) - - - -
dUltra-coding (ours) 81.52 23.54 (8.40) 35.64 33.39 (6.72) 35.85 24.29 (7.23) 37.04 19.78 (7.29)
   w/o planner 75.89 15.27 (12.84) 32.60 22.49 (9.89) 23.78 16.18 (10.19) 20.40 12.18 (11.28)
Table 1: Main results comparing dUltra with baseline methods across mathematical reasoning and code generation tasks. The generation length is 256 and the block size is 32 for all results in this table. We report accuracy, NFE (number of function evaluations), and T/F (tokens per forward pass) for each task. Bold indicates the best result and indicates the second-best result for each column. The “w/o planner” rows show results using Fast-dLLM sampling with the base model’s logits (without the learned unmasking planner head), serving as an ablation baseline. dUltra-math is trained on GSM8K training dataset while dUltra-coding uses dUltra-math as the base checkpoint and is trained on the APPS dataset.